Suppose | a b c |
| d e f | =8
| g h i |

Find value of | (g+2a) (h+2b) (i + 2c) |
| 3a 3b 3c |
| 2d 2e 2f |

I got 48 as my answer. please confirm if correct.
1
Expert's answer
2020-04-02T13:42:05-0400

"\begin{vmatrix}
a & b & c \\
d & e & f \\
g & h & i
\end{vmatrix}=8"


"\begin{vmatrix}
g+2a & h+2b & i+2c \\
3a & 3b & 3c \\
2d & 2e & 2f
\end{vmatrix}=
3\times 2\times 
\begin{vmatrix}
g+2a & h+2b & i+2c \\
a & b & c \\
d & e & f
\end{vmatrix}="

"= 6\times 
\begin{vmatrix}
g& h& i\\
a & b & c \\
d & e & f
\end{vmatrix}+6\times 
\begin{vmatrix}2a & 
2b &2c \\
a & b & c \\
d & e & f
\end{vmatrix}=-6\times \begin{vmatrix}
a & b & c \\
g & h & i \\
d & e & f
\end{vmatrix}+6\times 0="

"=(-1)^2 \times 6\times \begin{vmatrix}
a & b & c \\

d & e & f
\\
g & h & i \\
\end{vmatrix}=6\times 8=48"

Answer: 48.